A company manufactures and sales x computers per month. The monthly cost and price demand equation are C(x) = 72000 + 60𝑥 Price-demand equation is 𝑝(𝑥) = 200 − 𝑥 30 a. Find the maximum revenue. b. Find the maximum profit and the production level that will realize the maximum profit.  c. The price the company should sell for each computer. 


1
Expert's answer
2021-07-13T16:14:20-0400

a. Revenue is


"R(x)=p(x)x=(200-\dfrac{x}{30})x=200x-\dfrac{x^2}{30}, x\geq0"

"R'(x)=(200x-\dfrac{x^2}{30})'=200-\dfrac{x}{15}"

Critical number(s)


"R'(x)=0=>200-\dfrac{x}{15}=0=>x=3000"

If "0<x<3000, R'(x)>0, R(x)" increases.

If "x>3000, R'(x)<0, R(x)" decreases.

The function "R(x)" has a local maximum at "x=3000."

Since the function "R(x)" has the only extremum for "x\geq 0," then the function "R(x)" has the absolute maximum at "x=3000" for "x\geq0."


"R(3000)=200(3000)-\dfrac{(3000)^2}{30}=\$300000"

b.  Profit is


"P(x)=R(x)-C(x)=200x-\dfrac{x^2}{30}-(72000+60x)"

"=200x-\dfrac{x^2}{30}-(72000+60x)"

"=120x-\dfrac{x^2}{30}-72000"

"P'(x)=(120x-\dfrac{x^2}{30}-72000)'=120-\dfrac{x}{15}"

Critical number(s)


"P'(x)=0=>120-\dfrac{x}{15}=0=>x=1800"

If "0<x<1800, P'(x)>0, P(x)" increases.

If "x>1800, P'(x)<0, P(x)" decreases.

The function "P(x)" has a local maximum at "x=1800."

Since the function "P(x)" has the only extremum for "x\geq 0," then the function "P(x)" has the absolute maximum at "x=1800" for "x\geq0."


"P(1800)=120(1800)-\dfrac{(1800)^2}{30}=\$108000"



c.


"p(1800)=200-\dfrac{1800}{30}=\$140"